OpenStudy (anonymous):

Will FAN AND MEDAL!!! Will you plz help me find the answer for B, AND C E. Using the two points that your line of best fit goes through, complete the following: (1,50) and (7,100) a. What is the slope of the line of best fit? (1 point) 25/3 b. What is the equation of the line in point-slope form? (1 point) c. What is the equation of the line in slope-intercept form? (1 point)

OpenStudy (4everaddicted2anime):

I'm not sure how to do letter b, but I can do letter c. (1,50)(7,100) M=50/6 Y=(50/6)x + b 50=(50/6)(1) + b 50=(50/6) + b 250/6=B Y=(50/6)x + (250/6)